# supplementary_materials

### v1

@@ Appendix G Linear Response
$\mathcal{L}_{s}[\phi,A_{c},A_{v}]=\frac{\rho_{G}}{2}\left(\partial_{\mu}\phi-\mathbf{Q}-(A_{c,\mu}-A_{v,\mu})\right)^{2}$ (53)
@@
$\mathcal{L}_{s}[A_{c},A_{v}]=\frac{\rho_{G}}{2}\left(A_{c,\mu}-A_{v,\mu}\right)^{2}$ (54)
@@
$\mathcal{L}_{c}[\phi,A_{c},A_{v}]=\frac{\rho_{G}}{2}\left[\left(\partial_{\mu}\phi_{+}-\mathbf{Q}-(A_{c,\mu}-A_{v,\mu})\right)^{2}+\left(\partial_{\mu}\phi_{-}+\mathbf{Q}-(A_{c,\mu}-A_{v,\mu})\right)^{2}\right]$ (55)
@@
$\mathcal{L}_{c}[A_{c},A_{v}]=\rho_{G}\left(A_{c,\mu}-A_{v,\mu}\right)^{2}$ (56)
@@
The arguments presented above apply directly to the PDW superconductor by simply replacing $A_{c}-A_{v}$ by $A_{c}+A_{v}$ since the Cooper pair is electrically charged. One crucial consequence is that if the two layers are not separately contacted, $A_{c}=A_{v}$, the PDW superconductor indeed exhibits superflow, whereas, the X-DW phases do not, since the phase fields ϕ±subscriptitalic-$\phi_{\pm}$ decouple from the gauge fields.

### v2

@@ VII Linear Response
$\mathcal{L}_{s}[\phi,A_{c},A_{v}]=\frac{\rho_{G}}{2}\left(\partial_{\mu}\phi-\mathbf{Q}-(A_{c,\mu}-A_{v,\mu})\right)^{2}$ (A.50)
@@
$\mathcal{L}_{s}[A_{c},A_{v}]=\frac{\rho_{G}}{2}\left(A_{c,\mu}-A_{v,\mu}\right)^{2}$ (A.51)
@@
$\mathcal{L}_{c}[\phi,A_{c},A_{v}]=\frac{\rho_{G}}{2}\left[\left(\partial_{\mu}\phi_{+}-\mathbf{Q}-(A_{c,\mu}-A_{v,\mu})\right)^{2}+\left(\partial_{\mu}\phi_{-}+\mathbf{Q}-(A_{c,\mu}-A_{v,\mu})\right)^{2}\right]$ (A.52)
@@
$\mathcal{L}_{c}[A_{c},A_{v}]=\rho_{G}\left(A_{c,\mu}-A_{v,\mu}\right)^{2}$ (A.53)
@@
The arguments presented above apply directly to the PDW superconductor by simply replacing $A_{c}-A_{v}$ by $A_{c}+A_{v}$ since the Cooper pair is electrically charged. One crucial consequence is that if the two layers are not separately contacted, $A_{c}=A_{v}$, the PDW superconductor indeed exhibits superflow, whereas, the X-DW phases do not, since the phase fields $\phi_{\pm}$ decouple from the gauge fields.
Though our mean-field analysis suggest the formation of collinear X-DW over spiral X-DW, it is instructive to determine whether this may be experimentally verified. The collinear and the spiral X-DW phases show the same qualitative behavior in drag transport. However, the presence of intra-layer charge density modulation in the case of the collinear X-DW can distinguish the two possibilities, for example by scanning tunneling microscopy.
Finally, we discuss the response of the system to an in-plane magnetic field. The orbital coupling of the magnetic field will cause the two-dimensional momenta of the electrons in the two layers to shift in opposite directions, explicitly breaking $C_{2z}$. (For magnetic field along $\hat{x}$, the vector potential in the Landau gauge, $\mathbf{A}\sim z\hat{y}$, which is opposite for the two layers located at $z\pm d/2$.)The equally susceptible wave-vectors for exciton condensation will then be $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$ for $\mathbf{Q}_{1}\neq\mathbf{Q}_{2}$. In this case, we expect the collinear X-DW to get modified to a quasi-periodic ordered state where the strength of the exciton order parameter varies with a wavevector $(\mathbf{Q}_{1}+\mathbf{Q}_{2})/2$ and the phase of the order parameter varies with $(\mathbf{Q}_{1}-\mathbf{Q}_{2})/2$.